OpenStudy (anonymous):

Sweet corn of a certain variety is known to produce individual ears of corn with a mean weight of 8 ounces. A farmer is testing a new fertilizer designed to produce larger ears of corn, as measure by their weight. He finds that 28 randomly-selected ears of corn grown with this fertilizer have a mean weight of 8.25 ounces and a standard deviation of 0.8 ounces. There are no outliers in the data. Do these samples provide convincing evidence at the .05 level that the fertilizer had a positive impact on the weight of the corn ears?

ganeshie8 (ganeshie8):

figure out null/alternative hypotheses first

OpenStudy (anonymous):

yeaah I have no idea how to do this

ganeshie8 (ganeshie8):

\(H_0 : \mu = 8\) \(H_A : \mu \gt 8\)

ganeshie8 (ganeshie8):

find the p value

ganeshie8 (ganeshie8):

\(\large P(8.25~ |~ H_0~is~true)\)

ganeshie8 (ganeshie8):

to find p value, u need "standard error" u knw how to find it ha ?

OpenStudy (anonymous):

I know that Ha is the alternative hypothesis is always written as an inequality but that is all

OpenStudy (anonymous):

and is *

ganeshie8 (ganeshie8):

standard error = \(\large \frac{s}{\sqrt{n}} = \frac{0.8}{\sqrt{28}} = 0.1512\)

ganeshie8 (ganeshie8):

sketch the normal curve with below parameters : \(\mu = 8\) \(SE = 0.1512\)

ganeshie8 (ganeshie8):